OpenStudy (anonymous):

1 over 5 (x + 3) = 1 over 6 (x + 7)

OpenStudy (anonymous):

\[1/5(x+3) = 1/6(x+7) or 1/[5(x+3)] = 1/[6(x+7)] ?\]

OpenStudy (anonymous):

1/5(x+3)=1/6(x+7)

OpenStudy (anonymous):

'k multiply through by 30: 30*1/5(x+3)=30*1/6(x+7) 6(x+3) = 5(x+7) then distribute and solve

OpenStudy (anonymous):

\[\frac{1}{5 (3+x)}=\frac{1}{6 (7+x)}\]Move the RHS to the LHS, combine fractions, set the Numerator to zero and solve for x.\[\frac{27+x}{30 (3+x) (7+x)} \]\[\frac{27+x}{30 (3+x) (7+x)}=0 \]27x=0 x = -27

OpenStudy (anonymous):

6x+18 = 5x + 35

OpenStudy (anonymous):

mistake, 27 + x = 0.

OpenStudy (anonymous):

x = 35 - 18, whatever that is...

OpenStudy (anonymous):

according to the answer key, x = 17

OpenStudy (anonymous):

x = 17 is the solution to the following:\[\frac{x+3}{5}=\frac{x+7}{6} \]I interpreted the problem equation as follows:\[\frac{1}{5 (x+3)}=\frac{1}{6 (x+7)} \]